### v1

# Geometry of efficient weight vectors

Abstract
Pairwise comparison matrices and the weight vectors obtained from them are important concepts in multi-criteria decision making. A weight vector corresponding to a pairwise comparison matrix is called Pareto efficient if the approximation of the matrix elements by the weight ratios cannot be improved for any element of the matrix without worsening it for an other element. The aim of this paper is to show the geometrical properties of the set of Pareto efficient weight vectors for $4\times 4$ pairwise comparison matrices.
* Corresponding author, $1$ Kende u. 13-17., Budapest, Hungary;
Email: szadoczki.zsombor@sztaki.hun-ren.hu
a Research Group of Operations Research and Decision Systems,
Research Laboratory on Engineering & Management Intelligence
HUN-REN Institute for Computer Science and Control (SZTAKI), $1$ Kende u. 13-17., Budapest, Hungary;
Email: szadoczki.zsombor@sztaki.hun-ren.hu, bozoki.sandor@sztaki.hun-ren.hu
b Department of Operations Research and Actuarial Sciences
Corvinus University of Budapest, 1093 Fővám tér 8., Budapest, Hungary
Email: kristof.abele-nagy@uni-corvinus.hu
Keywords: Pairwise comparison, Pareto optimality, Geometry, Multi-criteria decision making, Analytic Hierarchy Process
1 Introduction
Pairwise comparisons are popular in many fields, such as decision making, psychology, and sports (Thurstone,, 1927; Triantaphyllou,, 2000; Csató,, 2021). Pairwise comparison matrices, originally developed for the Analytic Hierarchy Process (Saaty,, 1977, 1980), are important tools of multi-criteria decision making. When faced with a decision problem that involves multiple criteria, it is often not feasible to give the importance weights of the criteria directly. Instead, comparing the importance of the criteria pairwise in a ratio form by recording the answers to the questions ‘How many times is Criterion $i$ more important than Criterion $j$?’ is much more feasible. These answers can be arranged in a matrix, which is called a pairwise comparison matrix (PCM). The next step is to obtain a weight vector from the pairwise comparison matrix. This weight vector represents the estimation of the real underlying preferences of the decision maker. Instead of criteria, this approach can be applied in the same way to the performance of alternatives with respect to a single criterion. In this paper, the latter narrative is followed.
A PCM provided by a decision maker usually contains some contradictions, a certain level of inconsistency can be observed in the data. This can be measured by several inconsistency indices proposed in the literature (Brunelli,, 2018).
It is also worth mentioning that to estimate the decision maker’s underlying preferences (calculate a weight vector), not all of the pairwise comparisons are necessary. When there are missing elements in a PCM, then it is called an incomplete pairwise comparison matrix (Harker,, 1987). The structure of the known pairwise comparisons is often represented by undirected graphs (Gass,, 1998).
Since the matrix elements are the judgments of the decision maker, the weight ratios (regardless of how the weight vector is obtained), which are the approximations of the judgments, should be as close as possible to the matrix elements. If no approximation can be improved by changing the weights for any matrix element without making the approximation worse in some other element (thus, there is not another weight vector that (Pareto) dominates this one), the weight vector is called (Pareto) efficient. This is a natural requirement of a weight vector, since if the approximation of judgments can be improved in a trivial way, that improvement should be made.
Unfortunately, even popular weight vector calculation methods can produce inefficient weight vectors (Bozóki,, 2014). Thus, describing the set of efficient weight vectors is an important task. The efficiency of weight vectors has recently gained more attention resulting in several new findings and studies in the related literature. However, almost all of those are using the algebraic approach, while this paper aims to contribute to the complementary geometric view.
Blanquero et al., (2006) besides dealing with the efficiency of some of the most popular weight calculation techniques, connected the problem of efficient weight vectors to a directed graph, which was used widely later on.
Ábele-Nagy and Bozóki, (2016) and Ábele-Nagy et al., (2018) examined the popular eigenvector weight calculation method, and found that it is efficient for simple perturbed and double perturbed PCMs, i.e., for those ones, which only differ from a consistent matrix by one or two element(s) (and their reciprocal values), respectively. This result was extended to a certain class of triple perturbed PCMs—which differ from the consistent case by three elements (and their reciprocals)—by Fernandes and Furtado, (2022). Furtado and Johnson, 2024b showed different classes of PCMs for which the eigenvector method is (in)efficient.
The efficiency of so-called block perturbed PCMs was studied by Furtado and Johnson, 2024a . All efficient weight vectors of simple da Cruz et al., (2021) and double Furtado, (2023) perturbed PCMs were also described in an algebraic way in the literature.
Conde and de la Paz Rivera Pérez, (2010) used linear optimization problems to derive efficient weight vectors and determined a consistency index with an associated weakly efficient weight vector, while Bozóki and Fülöp, (2018) also developed LP models and algorithms to improve an inefficient weight vector, and proved that the eigenvector method is weakly efficient.
The efficiency of weight vectors calculated from PCMs gained in a real-world problem was examined by Duleba and Moslem, (2019).
Furtado and Johnson, 2024e proved that the geometric mean of any collection of distinct columns of a PCM is an efficient weight vector, which was then extended by Furtado and Johnson, 2024c , who showed that any Hadamard weighted geometric mean of the columns of a PCM is efficient. Furtado and Johnson, 2024d also examined the efficiency of the convex hull of the columns of certain PCMs.
Furtado and Johnson, 2024f deal with PCMs for which all efficient weight vectors provide the same ranking of alternatives, while Furtado and Johnson, (2025) provided an algebraic algorithm that determines inductively the complete set of efficient weight vectors for a PCM, as well as proved that the set of efficient weight vectors is piecewise linear connected.
Brunelli and Fedrizzi, (2024) formulated a property based on Pareto dominance related to the axiomatic systems used for the validation of inconsistency indices.
Szádoczki and Bozóki, (2024) examined the efficiency of weight vectors from a geometric viewpoint, and showed (mainly through examples) that the set of efficient weight vectors for a PCM with $3$ alternatives can be represented as a triangle (or a single point in the case of a consistent matrix), while for $4$ alternatives, when there are no consistent triads (triplets) among the alternatives, it can be visualized as a union of three tetrahedra. These tetrahedra are determined by the weight vectors calculated from those incomplete pairwise comparison matrices that can be obtained from the original PCM by deleting some elements, and can be represented with a line spanning tree graph. However, they did not characterize the possible cases, where the geometric properties of the set of efficient weight vectors differ, nor listed the possible cases. This paper aims to fill in this research gap, and determine the possible type of geometric characteristic of the set of efficient weight vectors, as well as show the underlying reasons, in the case of $4$ alternatives.
The rest of the paper is organized as follows. Section 2 contains the preliminaries related to pairwise comparison matrices and efficient weight vectors used throughout the study. Section 3 presents the main results, the geometric characterization of the possible cases of the set of efficient weight vectors. Finally, Section 4 concludes and raises further research questions.
2 Preliminaries
Definition 1 (Pairwise comparison matrix (PCM) Saaty, (1977)).
An $\mathbf{A}=\left[a_{ij}\right]_{i,j=1,\ldots,n}$ square matrix of dimensions $n\times n$ is a pairwise comparison matrix (PCM) if the following properties hold for all $i,j=1,\dots,n$: $a_{ij}>0$ (positivity) and $a_{ij}=1/a_{ji}$ (reciprocity).
Note that the second property implies $a_{ii}=1$. One of the most important concept related to pairwise comparison matrices is the consistency.
Definition 2 (Consistency Saaty, (1977)).
An $\mathbf{A}=\left[a_{ij}\right]_{i,j=1,\ldots,n}$ PCM is called consistent if $a_{ij}a_{jk}=a_{ik}$ for all $i,j,k=1,\dots,n$. A PCM is called inconsistent if it is not consistent.
There are several methods to obtain a weight vector related to a PCM, among which some of the most popular ones are the right eigenvector method (Saaty,, 1977), the logarithmic least squares (geometric mean) method (Crawford and Williams,, 1985), the least squares method (Jensen,, 1984), and the method based on the enumeration of spanning trees (Tsyganok,, 2010; Lundy et al.,, 2017; Mazurek and Kułakowski,, 2022) (for further techniques and their analysis, see, for instance Golany and Kress, (1993); Choo and Wedley, (2004); Bajwa et al., (2008); Dijkstra, (2013)). The ratios of the weights are considered to be the approximations of the elements of the PCM. Different methods can produce different weight vectors. The exact method of obtaining the weight vector will not be important for our purposes.
Pareto efficiency or Pareto optimality (Ehrgott,, 2000, Chapter 2)(Steuer,, 1986, Chapter 6) is an important concept of multi-objective optimization as well as multi-criteria decision making. The Pareto efficiency, also simply called the efficiency of a weight vector corresponding to a PCM can be defined such that by changing the elements of the weight vector, the ratios of the weights cannot approximate any element of the PCM better, without making the approximation of another element worse. The formal definition can be written as follows.
Definition 3 (Efficient weight vector).
Let $\mathbf{A}=\left[a_{ij}\right]_{i,j=1,\ldots,n}$ be an $n\times n$ PCM and $\mathbf{w}=(w_{1},w_{2},\ldots,w_{n})^{T}$ be a positive weight vector. $\mathbf{w}$ is called efficient if no other positive weight vector $\mathbf{w^{\prime}}=(w^{\prime}_{1},w^{\prime}_{2},\ldots,w^{\prime}_{n})^{T}$ exists such that
$\left|a_{ij}-\frac{w^{\prime}_{i}}{w^{\prime}_{j}}\right|$ $\leq\left|a_{ij}-\frac{w_{i}}{w_{j}}\right|\qquad\text{ for all }1\leq i,j\leq n,$ (1)
$\left|a_{k{\ell}}-\frac{w^{\prime}_{k}}{w^{\prime}_{\ell}}\right|$ $<\left|a_{k{\ell}}-\frac{w_{k}}{w_{\ell}}\right|\qquad\text{ for some }1\leq k,\ell\leq n.$ (2)
A weight vector is called inefficient if it is not efficient.
From the above definition it follows that renormalization does not affect efficiency or inefficiency.
Remark 1.
Weight vector $\mathbf{w}$ is efficient if and only if $c\mathbf{w}$ is efficient for any $c>0$.
Since for a consistent PCM $\mathbf{A}=\left[a_{ij}\right]_{i,j=1,\ldots,n}$ there exists (Saaty, (1977)) a weight vector $\mathbf{w}=(w_{1},w_{2},\ldots,w_{n})^{T}$ such that $a_{ij}=w_{i}/w_{j}$, this $\mathbf{w}$ is the only efficient weight vector (aside from a positive scalar multiplication) in this case.
Blanquero et al., (2006) devised a necessary and sufficient condition to the efficiency of a weight vector, using the following definition.
Definition 4 (Blanquero-Carrizosa-Conde (BCC) directed graph).
[…]
  • the vertices $V=\{1,2,\ldots,n\}$ correspond to the alternatives,
[…]
The efficiency of a weight vector can then be written by the following necessary and sufficient condition.
Theorem 1 ((Blanquero et al.,, 2006, Corollary 10)).
Let $\mathbf{A}$ be an $n\times n$ PCM. A weight vector $\mathbf{w}$ is efficient if and only if the BCC graph $G=(V,\overrightarrow{E})_{\mathbf{A},\mathbf{w}}$ is a strongly connected digraph, that is, there exist directed paths from $i$ to $j$ and from $j$ to $i$ for all pairs of nodes $i,j$.
For any pairs of vertices of the BCC digraph, there is an arc in at least one of the two possible orientations, and there are arcs in both directions only if the comparison of the appropriate alternatives is estimated perfectly by the given weight vector. This latter case can only happen at the boundary of the set of efficient weight vectors (Szádoczki and Bozóki,, 2024), which is a closed set (Furtado and Johnson, 2024e, , Theorem 6), thus, the focus of this paper is the case when there is exactly one arc for each pair of vertices, i.e., the relevant BCC digraphs are tournaments.
Definition 5 (Tournament).
A tournament is a directed graph in which each pair of vertices is connected by an arc with exactly one of the two possible orientations.
Theorem 2 (Camion, (1959)).
Let $G$ be a tournament. Then, $G$ is strongly connected if and only if $G$ has a Hamiltonian cycle (a cycle that visits each vertex exactly once).
Corollary 1.
A weight vector $\mathbf{w}$ associated with an $n\times n$ PCM $\mathbf{A}$ is efficient if and only if its BCC digraph $G=(V,\overrightarrow{E})_{\mathbf{A},\mathbf{w}}$ has a Hamiltonian cycle.
Remark 2.
For a BCC digraph with $n=3$, there are two possible Hamiltonian cycles (the $3$-cycle with $2$ different orientations), while for $n=4$ there are six possible Hamiltonian cycles (see the cycles in Figure 2).
As in Furtado and Johnson, (2025), the following notations will be used in the paper. Let us denote
  • the set of efficient weight vectors for the $n\times n$ PCM $\mathbf{A}$ by $\mathcal{E}\left(\mathbf{A}\right)$;
  • the vector obtained from vector $\mathbf{w}$ by deleting (retaining) the entry indexed by $i$ as $\mathbf{w}(i)$;
  • the submatrix of matrix $\mathbf{A}$ obtained by deleting (retaining) the column and row indexed by $i$ as $\mathbf{A}(i)$;
  • the set of $\mathbf{w}$ weight vectors for which $\mathbf{w}\in\mathcal{E}\left(\mathbf{A}\right)$, and $\mathbf{w}(i)\in\mathcal{E}\left(\mathbf{A}(i)\right)$ by $\mathcal{E}\left(\mathbf{A};i\right)$.
Theorem 3 ((Furtado and Johnson,, 2025, Corollaries 13 and 14)).
Let $\mathbf{A}$ be an $n\times n$ PCM, where $n\geq 4$. Then the following hold.
  1. 1.
    $\ \mathcal{E}\left(\mathbf{A}\right)=\bigcup_{i=1}^{n}\mathcal{E}\left(\mathbf{A};i\right)=\bigcup_{i=1,\ i\neq p}^{n}\mathcal{E}\left(\mathbf{A};i\right),$
    for any $p\in\{1,\ldots,n\}$.
  2. 2.
    For any $\mathbf{w}\in\mathcal{E}\left(\mathbf{A}\right)$ there are at least two submatrices of $\mathbf{A}$, $\mathbf{A}(i)$ and $\mathbf{A}(j)$ ($i\neq j$), such that $\mathbf{w}(i)\in\mathcal{E}\left(\mathbf{A}(i)\right)$ and $\mathbf{w}(j)\in\mathcal{E}\left(\mathbf{A}(j)\right).$
This means that any $\mathbf{w}$ weight vector that is efficient for PCM $\mathbf{A}$ must be efficient for at least two submatrices of $\mathbf{A}$, where exactly one row, the corresponding column, and the related entry of the weight vector are deleted. Furthermore, the set of efficient weight vectors ($\mathcal{E}\left(\mathbf{A}\right)$) can be determined by the union of any $n-1$ of the $n$ different sets of weight vectors that are efficient for $\mathbf{A}$ and for a given submatrix of $\mathbf{A}$ ($\mathbf{A}(i)$, $i=1,\ldots,n$).
Corollary 2.
Let $\mathbf{A}$ be an $n\times n$ PCM, where $n\geq 4$, and let $\mathbf{w}$ be a weight vector for which $\mathbf{w}\in\mathcal{E}\left(\mathbf{A};i\right)$. Then all the arcs of the BCC graph $G=(V,\overrightarrow{E})_{\mathbf{A},\mathbf{w}}$ that do not include the entry indexed $i$—denoted by $G(i)=\left(V(i),\overrightarrow{E}\right)_{\mathbf{A}(i),\mathbf{w}(i)}$—are determined by $\mathbf{w}(i)$ and $\mathbf{A}(i)$, and $G(i)$ must be strongly connected, thus must have a Hamiltonian cycle.
3 Results
The goal of this paper is to determine the possible type of geometric characteristic of the set of efficient weight vectors, and to show the reasons behind that for $4$ alternatives. However, the special cases, when there are consistent triads in the PCM, are not considered, for some geometric interpretation of those, see Szádoczki and Bozóki, (2024).
Assumption 1.
There is not a single consistent triad in the examined pairwise comparison matrix.
A general $4\times 4$ pairwise comparison matrix has the following form:
$\begin{pmatrix}1&a&b&c\\ 1/a&1&d&e\\ 1/b&1/d&1&f\\ 1/c&1/e&1/f&1\end{pmatrix},$
where $a,b,c,d,e,f>0$. From now on, we use these notations.
As the elements of a positive weight vector $\mathbf{w}$ are usually normalized to one, all of the possible vectors are in the unit $4$-simplex, which can be drawn in $3$ dimensions as a regular tetrahedron (denoted by $S_{4}$), using the following transformation.
$\mathbf{Tw}=\mathbf{\hat{w}},$
where
$\mathbf{T}=\begin{pmatrix}1&1&0&0\\ 1&0&1&0\\ 0&1&1&0\end{pmatrix}.$
Thus, a given $4$-dimensional weight vector $\mathbf{w}$ can be represented by $\mathbf{\hat{w}}=(x,y,z)^{T}$ in $S_{4}$. It is also true that the efficiency of weight vector $\mathbf{w}$ is determined by whether the related BCC graph $G=(V,\overrightarrow{E})_{\mathbf{A},\mathbf{w}}$ is strongly connected. Using the general $4\times 4$ PCM notations, the arcs of $G$ are determined by the following hyperplanes:
$h_{12}:\ w_{1}/w_{2}=a,$
$h_{13}:\ w_{1}/w_{3}=b,$
$\ldots$
$h_{34}:\ w_{3}/w_{4}=f.$
Figure 1 shows $S_{4}$ with an example plane of $h_{12}:$ $w_{1}/w_{2}=a=2$.
[…]
Figure 1: $S_{4}$ with an example plane of $h_{12}:$ $w_{1}/w_{2}=a=2$ and the related arcs of the BCC graph above and below the plane.
One can see that the possible weight vectors are inside $S_{4}$ (all the elements of the vectors are positive), while the vertices of $S_{4}$ ($V_{1},V_{2},V_{3}$ and $V_{4}$) correspond to the cases, where exactly one element is 1, and all the others are 0 (transformation matrix $\mathbf{T}$ is also applied here). The example plane ($h_{12}$) determines the arc of the $G$ BCC graph between alternatives $1$ and $2$, i.e., if $w_{1}/w_{2}>2$ holds, then there is an arc from $1$ to $2$, and if $w_{1}/w_{2}<2$, then the arc is from $2$ to $1$. On the plane itself both directions hold. It is also notable that the plane splits the $V_{1}V_{2}$ edge of $S_{4}$ according to a $1:2$ ratio.
For every element of the PCM (aside from the ones coming from the reciprocity property) there is such a plane that determines one of the arcs of the BCC graph. According to Theorem 1, if the gained BCC graph is strongly connected, then the related weight vector is efficient. The question is how can we characterize the set of efficient weight vectors in a geometric way. How many different types of sets are possible from a geometric view point, and what is the determining factor behind that?
To answer these questions, first the pairwise comparison matrices that should be considered are narrowed down.
If all of the possible vectors and planes are multiplied in a given dimension by a constant, then geometrically that is only a stretching, the relation of any two vectors are not modified. This observation is formalized for the examined problem in Proposition 1.
Proposition 1.
It is sufficient to consider matrices in the form of
$\begin{pmatrix}1&1&b^{\prime}&c^{\prime}\\ 1&1&1&e^{\prime}\\ 1/b^{\prime}&1&1&1\\ 1/c^{\prime}&1/e^{\prime}&1&1\end{pmatrix}.$
Proof.
Let us consider
$\mathbf{A}=\begin{pmatrix}1&a&b&c\\ 1/a&1&d&e\\ 1/b&1/d&1&f\\ 1/c&1/e&1/f&1\end{pmatrix},\mathbf{w}=\begin{pmatrix}w_{1}\\ w_{2}\\ w_{3}\\ w_{4}\end{pmatrix},$
where $\mathbf{w}$ is the weight vector obtained from $\mathbf{A}$ with any chosen weight vector calculation method. Let us multiply the weight of the first alternative with $1/adf$ (a shorter notation for $1/(adf)$). This is akin to a measurement unit change. This will result in a corresponding change in $\mathbf{A}$. We get
$\begin{pmatrix}1&1/df&b/adf&c/adf\\ df&1&d&e\\ adf/b&1/d&1&f\\ adf/c&1/e&1/f&1\end{pmatrix},\begin{pmatrix}w_{1}/adf\\ w_{2}\\ w_{3}\\ w_{4}\end{pmatrix}.$
Now let us multiply the second alternative with $1/df$. We get
$\begin{pmatrix}1&1&b/adf&c/adf\\ 1&1&1/f&e/df\\ adf/b&f&1&f\\ adf/c&df/e&1/f&1\end{pmatrix},\begin{pmatrix}w_{1}/adf\\ w_{2}/df\\ w_{3}\\ w_{4}\end{pmatrix}.$
Now let us multiply the third alternative with $1/f$. We get
$\mathbf{A}^{\prime}=\begin{pmatrix}1&1&b/ad&c/adf\\ 1&1&1&e/df\\ ad/b&1&1&1\\ adf/c&df/e&1&1\end{pmatrix},\mathbf{w}^{\prime}=\begin{pmatrix}w_{1}/adf\\ w_{2}/df\\ w_{3}/f\\ w_{4}\end{pmatrix}.$
Thus, with $b^{\prime}=b/ad,c^{\prime}=c/adf,e^{\prime}=e/df$ the matrix has the required form. ∎
Note that after the transformation $\sum w_{i}=1$ no longer holds. This means that instead of $S_{4}$, the weight vectors should be illustrated in a tetrahedron, where the vertices are the following:
$(w_{1}=1/adf,w_{2}=0,w_{3}=0,w_{4}=0),$
$(w_{1}=0,w_{2}=1/df,w_{3}=0,w_{4}=0),$
$(w_{1}=0,w_{2}=0,w_{3}=1/f,w_{4}=0),$
$(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1).$
However, the geometric nature of the set of efficient weight vectors will be the same as the one for the matrices that follow the form presented in Proposition 1 in $S_{4}$. Thus, we will only focus on those to determine the different types of geometric characteristic of sets of efficient weight vectors.
It is also true that we can get the same geometric object with reversed dimensions (looking at it from different angles). This is formalized for pairwise comparison matrices in Proposition 2.
Proposition 2.
Any $4\times 4$ pairwise comparison matrix
$\mathbf{A}=\begin{pmatrix}1&a&b&c\\ 1/a&1&d&e\\ 1/b&1/d&1&f\\ 1/c&1/e&1/f&1\end{pmatrix}$
can be transformed with appropriate reindexing of alternatives to the following form:
$\mathbf{A}^{\prime\prime}=\begin{pmatrix}1&a^{\prime\prime}&b^{\prime\prime}&c^{\prime\prime}\\ 1/a^{\prime\prime}&1&d^{\prime\prime}&e^{\prime\prime}\\ 1/b^{\prime\prime}&1/d^{\prime\prime}&1&f^{\prime\prime}\\ 1/c^{\prime\prime}&1/e^{\prime\prime}&1/f^{\prime\prime}&1\end{pmatrix},$
where $\mathbf{A}^{\prime\prime}$ falls into one of the following cases:
Case 1: $a^{\prime\prime}d^{\prime\prime}<b^{\prime\prime}$, $d^{\prime\prime}f^{\prime\prime}<e^{\prime\prime}$, $b^{\prime\prime}f^{\prime\prime}<c^{\prime\prime}$, $a^{\prime\prime}e^{\prime\prime}<c^{\prime\prime}$.
Case 2: $a^{\prime\prime}d^{\prime\prime}<b^{\prime\prime}$, $d^{\prime\prime}f^{\prime\prime}<e^{\prime\prime}$, $b^{\prime\prime}f^{\prime\prime}<c^{\prime\prime}$, $a^{\prime\prime}e^{\prime\prime}>c^{\prime\prime}$.
Proof.
Let us consider $\mathbf{A}$ and let us swap the index of the first two alternatives. This will result in the following pairwise comparison matrix:
$\mathbf{A_{21}}=\begin{pmatrix}1&1/a&d&e\\ a&1&b&c\\ 1/d&1/b&1&f\\ 1/e&1/c&1/f&1\end{pmatrix}$
Let us assume that in $\mathbf{A}$
  • for triad (1,2,3) $ad>b$,
  • for triad (2,3,4) $df>e$,
  • for triad (1,3,4) $bf>c$,
  • for triad (1,2,4) $ae>c$.
Then, in $\mathbf{A_{21}}$
  • for triad (1,2,3) $(1/a)\cdot b<d$,
  • for triad (2,3,4) $bf>c$,
  • for triad (1,3,4) $df>e$,
  • for triad (1,2,4) $(1/a)\cdot c<e$.
This transformation reversed the direction of the relation in those triads where both the reindexed alternatives were involved, and left them in place where only one of them was involved. With the same reasoning, a further swap of the indices of alternatives 3 and 4 will result in the form in Case 1. This is true if any pair of triads have their relation in a $>$ direction, swapping their indices will result in the relations corresponding to those triads to become $<$. Thus, if exactly zero, two or four of the triads have their relation in a $>$ direction, performing zero, one, or two index swappings of the appropriate alternatives will result in the form of Case 1.
On the other hand, if one or three triads have their relation in a $>$ direction, performing an appropriate number of index swaps will result in the form of Case 2. ∎
Corollary 3.
Any $4\times 4$ pairwise comparison matrix can be transformed with appropriate reindexing and multiplication of weights of alternatives to one of the following forms:
$\begin{pmatrix}1&1&b^{\prime\prime}&c^{\prime\prime}\\ 1&1&1&e^{\prime\prime}\\ 1/b^{\prime\prime}&1&1&1\\ 1/c^{\prime\prime}&1/e^{\prime\prime}&1&1\end{pmatrix},$
Case 1: $1\cdot 1<b^{\prime\prime}$, $1\cdot 1<e^{\prime\prime}$, $b^{\prime\prime}\cdot 1<c^{\prime\prime}$, $1\cdot e^{\prime\prime}<c^{\prime\prime}$.
Case 2: $1\cdot 1<b^{\prime\prime}$, $1\cdot 1<e^{\prime\prime}$, $b^{\prime\prime}\cdot 1<c^{\prime\prime}$, $1\cdot e^{\prime\prime}>c^{\prime\prime}$.
Proof.
Apply Proposition 2 to the pairwise comparison matrix in Proposition 1. ∎
Corollary 3 means that it is enough to examine these two cases, and show the geometric nature of the set of efficient weight vectors, and the directions of the arcs in the strongly connected BCC graphs.
First, let us consider Case 1. Szádoczki and Bozóki, (2024) showed that the set of efficient weight vectors for $4$ alternatives geometrically consists of the union of three tetrahedra (denoted by $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$) associated with the labeled Hamiltonian cycles of BCC graphs presented in Figure 2, where for a given PCM always only one of the two directions is possible for each cycle.
[…]
Figure 2: The labeled Hamiltonian cycles of BCC graphs associated with tetrahedra $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$. Each tetrahedron is related to the cycles below them, from which always only one is possible for a given PCM.
As shown in Figure 1 there is a cutting plane for every element of the pairwise comparison matrix (aside from the elements coming from the reciprocity property). However, each of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$ is determined by exactly four planes. Figures 3, 4, and 5 show $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$, respectively, with the planes determining them, while Figure 6 present the set of efficient weight vectors (the union of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$) for the following illustrative example matrix ($b^{\prime\prime}=2$, $c^{\prime\prime}=4$, $e^{\prime\prime}=3$):
$\mathbf{A^{\prime\prime}}=\begin{pmatrix}1&1&2&4\\ 1&1&1&3\\ 1/2&1&1&1\\ 1/4&1/3&1&1\end{pmatrix}.$
Note that in Figures 3, and 4 plane $h_{23}$ (denoted by dark green color) is barely visible from this angle, as it connects $V_{1}$, $V_{4}$ and a point of the $V_{2}V_{3}$ edge that happens to coincide with a point from the $V_{1}V_{4}$ edge, thus from this plane only the $V_{1}V_{4}$ edge can be observed. This is true for all the latter figures containing $h_{23}$ as well.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$1234
Figure 3: The first tetrahedron of efficient weight vectors ($\mathcal{T}_{1}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{14},h_{23}$, and $h_{34}$) for the Case 1 example.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$1324
Figure 4: The second tetrahedron of efficient weight vectors ($\mathcal{T}_{2}$), the associated Hamiltonian cycle, and the determining planes ($h_{13},h_{14},h_{23}$, and $h_{24}$) for the Case 1 example.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$1243
Figure 5: The third tetrahedron of efficient weight vectors ($\mathcal{T}_{3}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{13},h_{24}$, and $h_{34}$) for the Case 1 example.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$123413241243
Figure 6: The set of efficient weight vectors for the Case 1 example (the union of $\mathcal{T}_{1},\mathcal{T}_{2}$ and $\mathcal{T}_{3}$).
As for Case 2, we consider the following illustrative example matrix ($b^{\prime\prime}=2$, $c^{\prime\prime}=3$, $e^{\prime\prime}=4$):
$\mathbf{A^{\prime\prime}}=\begin{pmatrix}1&1&2&3\\ 1&1&1&4\\ 1/2&1&1&1\\ 1/3&1/4&1&1\end{pmatrix}.$
Figures 7, 8, and 9 show $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$, respectively, with the planes determining them, while Figure 10 present the set of efficient weight vectors (the union of $\mathcal{T}_{1},\mathcal{T}_{2},$ and $\mathcal{T}_{3}$) for this example.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$1234
Figure 7: The first tetrahedron of efficient weight vectors ($\mathcal{T}_{1}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{14},h_{23}$, and $h_{34}$) for the Case 2 example.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$1324
Figure 8: The second tetrahedron of efficient weight vectors ($\mathcal{T}_{2}$), the associated Hamiltonian cycle, and the determining planes ($h_{13},h_{14},h_{23}$, and $h_{24}$) for the Case 2 example.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$1243
Figure 9: The third tetrahedron of efficient weight vectors ($\mathcal{T}_{3}$), the associated Hamiltonian cycle, and the determining planes ($h_{12},h_{13},h_{24}$, and $h_{34}$) for the Case 2 example.
$V_{1}=(w_{1}=1,w_{2}=0,w_{3}=0,w_{4}=0)$$V_{2}=(w_{1}=0,w_{2}=1,w_{3}=0,w_{4}=0)$$V_{3}=(w_{1}=0,w_{2}=0,$$w_{3}=1,w_{4}=0)$$V_{4}=(w_{1}=0,w_{2}=0,w_{3}=0,w_{4}=1)$$x$$y$$z$123413241243
Figure 10: The set of efficient weight vectors for the Case 2 example (the union of $\mathcal{T}_{1},\mathcal{T}_{2}$ and $\mathcal{T}_{3}$).
4 Conclusion and further research
The geometry of efficient weight vectors is relevant for larger pairwise comparison matrices as well, although visualization has already achieved its limit at $n=4$. The pattern shown above suggests that each of the $\frac{1}{2}(n-1)!$ Hamiltonian cycles of the BCC graph corresponds to a simplex, determined by $n$ spanning trees’ weight vectors. These simplices form building blocks of the set of efficient weight vectors. Within each simplex, the weight vectors associated to the paths of length $n-1$ form a local basis. Every efficient weight vector can be written as a convex combination of $n$ spanning trees’ weight vectors, belonging to the same Hamiltonian cycle. Note that all the other spanning trees’ (different from a path of length $n-1$) weight vectors can also be written in that way. Generalizations of Propositions 1 and 2 are also necessary in future research.
Acknowledgements
The research was supported by the National Research, Development and Innovation Office under Grants FK 145838 and TKP2021-NKTA-01 NRDIO.